# A Simple and Efficient Joint Measurement Strategy for Estimating Fermionic Observables and Hamiltonians

### v1

@@ Abstract
We propose a simple scheme to estimate fermionic observables and Hamiltonians relevant in quantum chemistry and correlated fermionic systems. Our approach is based on implementing a measurement that jointly measures noisy versions of any product of two or four Majorana operators in an $N$ mode fermionic system. To realize our measurement we use: (i) a randomization over a set of unitaries that realize products of Majorana fermion operators; (ii) a unitary, sampled at random from a constant-size set of suitably chosen fermionic Gaussian unitaries; (iii) a measurement of fermionic occupation numbers; (iv) suitable post-processing. Our scheme can estimate expectation values of all quadratic and quartic Majorana monomials to $\epsilon$ precision using $\mathcal{O}(N\log(N)/\epsilon^{2})$ and $\mathcal{O}(N^{2}\log(N)/\epsilon^{2})$ measurement rounds respectively, matching the performance offered by fermionic shadow tomography [1, 2]. In certain settings, such as a rectangular lattice of qubits which encode an $N$ mode fermionic system via the Jordan-Wigner transformation, our scheme can be implemented in circuit depth $\mathcal{O}(N^{1/2})$ with $\mathcal{O}(N^{3/2})$ two-qubit gates, offering an improvement over fermionic and matchgate classical shadows that require depth $\mathcal{O}(N)$ and $\mathcal{O}(N^{2})$ two-qubit gates. We also benchmark our method on molecular Hamiltonians and observe performances comparable to those offered by fermionic classical shadows.
@@ I Introduction
One of the most promising tasks in which quantum computers can offer practical speedups over their classical counterparts is the problem of determining low energy states of molecular Hamiltonians in quantum chemistry (see [3] for a review). While the ideal approach would involve simulating the quantum system on a (yet to be realized) universal quantum computer, a reasonable goal in the meantime is to find problems where smaller sized quantum devices can outperform current classical approaches. To this end, much effort has been spent developing algorithms that combine both quantum and classical techniques to achieve speedups [4, 5, 6, 7], e.g., for the simulation of time dynamics in quantum chemistry systems [8, 9]. A large class of these algorithms require an approximation of the ground state energy of a Hamiltonian, which involves estimating the Hamiltonian on a quantum device and optimizing the prepared state by classical methods.
@@
As the system size increases, one of the main performance bottlenecks is the inability to simultaneously measure the non-commuting observables which appear in the Hamiltonian’s decomposition. Non-commutativity, more generally, hinders the ability to efficiently learn many properties of a quantum system [10, 11, 12, 13, 14, 15]. To reduce the measurement cost (and the number of state preparations) several strategies have been developed. A significant body of work attempts to efficiently find minimal groupings of the measurements into sub-classes of commuting observables [16, 17, 18, 19, 20, 21, 22, 23]. Another approach, known as classical shadows [24], is based on a randomized measurement strategy that yields a classical approximation of the quantum state. The resulting approximation can be used to simultaneously estimate non-commuting observables in both many-body qubit [24, 25, 26] and fermionic systems [1, 2, 27, 28], with state-of-the-art sample complexity scalings.
@@
In this work we consider an alternative approach for estimating non-commuting fermionic observables by implementing a simple and efficient joint measurement, in analogy to the strategy for multi-qubit Pauli observables [29]. In particular, we construct a joint measurement of a modified (noisy) collection of Majorana monomials, whose outcome statistics can be reproduced from the (parent) measurement together with an efficient classical post-processing [30, 31, 32]. Our measurement scheme for an $N$ mode fermionic system involves sampling from two distinct subsets of unitaries followed by a measurement of the occupation numbers and a simple classical post-processing. The first subset of unitaries realizes products of Majorana fermion operators. The second subset (consisting of several fermionic Gaussian unitaries) rotates disjoint blocks of the Majorana operators into balanced superpositions. For the case of quadratic and quartic monomials we show that sets of two or nine fermionic Gaussian unitaries are sufficient, respectively, to jointly measure all noisy versions of the desired observables. We also describe a tailored joint measurement scheme for estimating energies of electronic structure Hamiltonians, where four fermionic Gaussian unitaries in the second subset are sufficient.
@@
Our joint measurement strategy estimates expectation values of quadratic and quartic Majorana monomials to $\epsilon$ precision with $\mathcal{O}(N\log(N)/\epsilon^{2})$ and $\mathcal{O}(N^{2}\log(N)/\epsilon^{2})$ measurement rounds, respectively, and provides the same performance guarantees as fermionic (matchgate) classical shadows [1, 2]. Under the Jordan-Wigner transformation, in the setting of a rectangular lattice of qubits, the measurement circuit can be achieved with depth $\mathcal{O}(N^{1/2})$ and using $\mathcal{O}(N^{3/2})$ two-qubit gates. Fermionic and matchgate classical shadow circuits, on the other hand, require depth $\mathcal{O}(N)$ and $\mathcal{O}(N^{2})$ two-qubit gates, as well as requiring a randomization over a large collection of unitaries [1, 2]. We benchmark our strategy with several popular molecular Hamiltonians and find that the sample complexity is comparable to other state-of-the-art schemes [24, 1, 33, 25].
@@
This work is accompanied by a complementary paper [34], which studies joint measurability of Majorana fermion operators from a more formal perspective, in particular by showing (asymptotic) optimality of our joint measurement scheme and establishing connections to graph and design theory, as well as topics in mathematical physics such as the SYK model.
@@
The rest of the paper is organized as follows. First, in Section II we introduce the necessary concepts regarding fermionic systems and quantum measurements. In Section III we present the general protocol for estimating expectation values of products of Majorana operators. In Section IV we provide the general structure of fermionic Gaussian unitaries that guarantee a good scaling of the variance of the estimators for Majorana pairs and quadruples. In Section V we discuss how to practically implement our scheme on a quantum computer under different fermion-to-qubit mappings and in different qubit layouts. In Section VI we specialize our scheme to quantum chemistry Hamiltonians and present its optimized implementation for a 2D qubit layout. Finally, we conclude the paper with a short discussion and list of open problems.
II Setting and notation
@@
It is convenient to introduce Majorana fermion operators $\gamma_{1},\gamma_{2},\ldots,\gamma_{2N}$, which are fermionic analogues of quadratures known from quantum optics and defined as $\gamma_{2i-1}=a_{i}+a^{\dagger}_{i}$ and $\gamma_{2i}=\mathrm{i}(a^{\dagger}_{i}-a_{i})$. Majorana operators are Hermitian and satisfy the anticommutation relations $\{{\gamma_{i},\gamma_{j}\}}=\delta_{ij}\mathbb{I}$. In this work we are interested in estimating expectation values of (Hermitian) products of Majorana operators, with particular emphasis of monomials of degree two and four (pairs and quadruples). For even-sized subset $A\subset[2N]$ (using the convention $[k]=\{{1,\ldots,k\}}$) we define: $\gamma_{A}\coloneqq\mathrm{i}^{|A|/2}\prod_{i\in A}\gamma_{i}$, where $|A|$ denotes the size of subset $A$. We assume the convention that terms in the product are arranged in a non-decreasing manner, for example $\gamma_{1,3}=i\gamma_{1}\gamma_{3}$ and $\gamma_{1,3,6,11}=-\gamma_{1}\gamma_{3}\gamma_{6}\gamma_{11}$. For any $A,B\subset[2N]$, Majorana monomials satisfy the commutation relation $\gamma_{A}\gamma_{B}=(-1)^{|A|\cdot|B|+|A\cap B|}\gamma_{B}\gamma_{A}$, which makes it impossible to measure all observables simultaneously [35].
@@
Figure 1: Two exemplary realizations of the Jordan-Wigner encoding in systems of qubits that exhibit different spatial locality: a) 1D array and b) $(L+1)\times L/2$ rectangular qubit array (for $L=4$). The individual qubits are indicated by consecutive pairs within the blue rounded rectangles labelled by $Q_{1},Q_{2},\ldots$, each enclosing two Majorana mode operators indicated by black dots. Importantly, locality in Majorana modes (e.g. coupling $\mathrm{exp}(\mathrm{i}\varphi\gamma_{i,j})$ for neighbouring Majorana modes $i,j$) does not in general correspond to locality with respect to the structure of the qubit lattice.
It is often convenient to represent a fermionic system as a system of qubits via a fermion-to-qubit mapping. A transformation of this type maps the $2N$ Majorana operators to multi-qubit Pauli operators while preserving their anticommutation relations. One such mapping is the Jordan-Wigner (JW) transformation [39], in which the Majorana operators of mode $j\in[N]$ are represented as the $N$-qubit Pauli operators, $\gamma_{2j-1}=Z^{\otimes j-1}\otimes X\otimes\mathbb{I}_{2}^{\otimes n-j}$ and $\gamma_{2j}=Z^{\otimes j-1}\otimes Y\otimes\mathbb{I}_{2}^{\otimes n-j}$, where $X,Y$ and $Z$ are the qubit Pauli operators, and $\mathbb{I}_{2}$ the $2\times 2$ identity. While the Jordan-Wigner transformation can be defined irrespective of the physical arrangement of qubits, different layouts (and labellings of qubits) can have implementation benefits (see Fig. 1 for exemplary realizations of the Jordan-Wigner transformation in 1D and 2D layouts).
@@
with $\eta_{A}\in(0,1]$ the visibility and $e_{A}\in\{\pm 1\}$ the measurement outcome.
@@
of the expectation values $\operatorname{tr}(\gamma_{A}\rho)$ for an (unknown) state $\rho$. We can therefore estimate many non-commuting operators simultaneously, just as in the case of classical shadows [1, 2, 24]. The efficiency of the strategy for a set of observables $\mathcal{S}_{k}=\{\gamma_{A}:|A|=k\}$ is quantified by the sample complexity, i.e. the number of copies of the state $\rho$ needed to ensure, with probability at least $1-\delta$, that $|\operatorname{tr}(\gamma_{A}\rho)-\hat{\gamma}_{A,S}|<\epsilon$ for all $\gamma_{A}\in\mathcal{S}_{k}$, where $\hat{\gamma}_{A,S}=(1/S)\sum_{i=1}^{S}\hat{\gamma}^{(i)}_{A}$ is the empirical mean of the estimator $\hat{\gamma}_{A}$ from $S$ independent rounds. Since the estimator $\hat{\gamma}_{A}$ is binary and takes the values {±$\{\pm\eta_{A}^{-1}\}$, we can apply Hoeffding’s inequality to bound the probability for each observable [40]. Following simple reasoning based on the union bound (over different observables $\gamma_{A}$), we get
@@
where $|\mathcal{S}_{k}|=\binom{2n}{k}$ denotes the number of estimated observables and $\eta_{k}$ the visibility for which all observables in $\mathcal{S}_{k}$ can be jointly measured.
@@
III A General Joint Measurement Strategy
@@
Figure 2: Circuit implementation of the joint measurement scheme described in (i)–(iv) in Sec. III for an $N$ mode fermionic state. In step (i) the state evolves under the action of a Majorana monomial $\gamma_{X}^{\dagger}$, sampled uniformly at random from the set of $2^{2N}$ Majorana monomials. In step (ii) a fermionic Gaussian unitary $U^{\dagger}$ is applied. In step (iii) the $N$ projective measurements of the disjoint Majorana pairs are performed, with outcomes $q_{1},\ldots,q_{N}$. Finally, for a given $A\subset[2N]$, one outputs $e_{A}$ according to step (iv).
@@
    Output $e_{A}=s_{AB}(-1)^{|A\cap X|}\prod_{j\in B^{\prime}}q_{i}$.
@@
Proof.
We first note that
where we have used (3) and the commutation relation of the operators $\gamma_{X}$ and $\gamma_{Y}$. In order to prove the claim in the proposition it suffices to show that for $x=\pm 1$ we have $\mathrm{Pr}_{X,\bm{q}}(e_{A}=x)=\operatorname{tr}(\rho\frac{1}{2}(\mathbb{I}+x\cdot\eta_{A}\gamma_{A}))$, where $X$ and $\bm{q}$ are random variables on which $e_{A}$ depends. Using the definition of $e_{A}$ in (iv) and the fact that the distribution of $X$ is uniform we have
$\frac{s_{AB}}{2^{2N+1}}\sum_{\begin{subarray}{c}X,Y\subseteq[2N]\\ |Y|=|B|\end{subarray}}\operatorname{tr}\left(\rho(-1)^{|A\cap X|+|X\cap Y|}\gamma_{Y}\det(R_{Y,B})\right)$ (12)
$=\frac{1}{2}\eta_{A}\operatorname{tr}(\rho\gamma_{A})\ ,$ (13)
where we have used that for $|A|,|B|$–even:
Note that steps (i)–(iii) above are independent of the observable $\gamma_{A}$ whose noisy variant we aim to measure. In step (iv) we conduct an efficient post-processing that depends only on $A$ and $B^{\prime}$ (which can depend on $A$) and a randomly chosen subset $X$ of $[2N]$. For this reason the POVM realized in steps (i)–(iii) is a parent POVM for noisy versions of the observables $\gamma_{A}$ with visibility parameter $\eta_{A}$, for all subsets $A\subset[2N]$ of even size. The outcomes can therefore be used to estimate, simultaneously, all non-commuting observables via the estimators $\hat{\gamma}_{A}$ from Eq. (5), since $\eta_{A}^{-1}\mathbb{E}_{X,\bm{q}}[e_{A}]=\operatorname{tr}(\rho\gamma_{A})$. Next, to minimise the sample complexity we construct unitaries which aim to maximise the visibility of the measurements.
IV Gaussian unitaries for measurements of Majorana pairs and quadruples
We propose a modification of the original scheme by introducing a preliminary step (o) in which every experimental run is preceded by sampling uniformly at random from a set of $K$ unitaries $U^{(r)}\in\mathrm{FLO}(N)$, $r=1,\ldots,K$. Then, we proceed with the usual steps (i)–(iii), as described in Sec. III, with $U^{(r)}$ implemented in (iii). Here we sketch the construction of the $K$ unitaries that ensures that the joint measurement from Prop. 1 (or rather its simple modification) gives visibilities $\eta_{2}=\Theta(1/\sqrt{N})$ and $\eta_{4}=\Theta(1/N)$, where $\eta_{k}$ denotes the maximum visibility for which all $k$–th degree Majorana monomials can be jointly measured by the parent. The details of the construction together with an extension to arbitrary constant-sized Majorana monomials are presented in [34].

### v2

@@ Abstract
We propose a simple scheme to estimate fermionic observables and Hamiltonians relevant in quantum chemistry and correlated fermionic systems. Our approach is based on implementing a measurement that jointly measures noisy versions of any product of two or four Majorana operators in an $N$ mode fermionic system. To realize our measurement we use: (i) a randomization over a set of unitaries that realize products of Majorana fermion operators; (ii) a unitary, sampled at random from a constant-size set of suitably chosen fermionic Gaussian unitaries; (iii) a measurement of fermionic occupation numbers; (iv) suitable post-processing. Our scheme can estimate expectation values of all quadratic and quartic Majorana monomials to $\epsilon$ precision using $\mathcal{O}(N\log(N)/\epsilon^{2})$ and $\mathcal{O}(N^{2}\log(N)/\epsilon^{2})$ measurement rounds respectively, matching the performance offered by fermionic shadow tomography [1, 2]. In certain settings, such as a rectangular lattice of qubits which encode an $N$ mode fermionic system via the Jordan-Wigner transformation, our scheme can be implemented in circuit depth $\mathcal{O}(N^{1/2})$ with $\mathcal{O}(N^{3/2})$ two-qubit gates, offering an improvement over fermionic and matchgate classical shadows that require depth $\mathcal{O}(N)$ and $\mathcal{O}(N^{2})$ two-qubit gates. By benchmarking our method on exemplary molecular Hamiltonians and observing performances comparable to fermionic classical shadows, we demonstrate a novel, competitive alternative to existing strategies.
@@ I Introduction
Some of the most promising tasks in which quantum computers can offer practical speedups over their classical counterparts can be found in the fields of quantum chemistry and simulation (see e.g. [3] for a review), where fermionic systems are of particular interest. While the ideal approach would involve simulating the quantum system on a (yet to be realized) universal quantum computer, a reasonable goal in the meantime is to find problems where smaller sized quantum devices can outperform current classical approaches, given the limitations on achievable gate depths and counts. To this end, much effort has been spent developing algorithms that combine both quantum and classical techniques to achieve speedups [4, 5, 6, 7]. These can target, e.g., the simulation of time dynamics in quantum chemistry systems [8, 9], as well as the problem of determining low energy states of molecular Hamiltonians. A large class of these algorithms, known as variational quantum algorithms [6], require an approximation of the ground state energy of a Hamiltonian, which involves estimating the Hamiltonian on a quantum device and optimizing the prepared state by classical methods.
@@
As the system size increases, one of the main performance bottlenecks in the aforementioned tasks is the inability to simultaneously measure the relevant non-commuting observables, e.g. those which appear in the Hamiltonian’s decomposition. Non-commutativity, more generally, hinders the ability to efficiently learn many properties of a quantum system [10, 11, 12, 13, 14, 15]. To reduce the measurement cost (and the number of state preparations) several strategies have been developed. A significant body of work attempts to efficiently find minimal groupings of the measurements into sub-classes of commuting observables [16, 17, 18, 19, 20, 21, 22, 23]. Another approach, known as classical shadows [24], is based on a randomized measurement strategy that yields a classical approximation of the quantum state. The resulting approximation can be used to simultaneously estimate non-commuting observables in both many-body qubit [24, 25, 26] and fermionic systems [1, 2, 27, 28], with state-of-the-art sample complexity scalings. A relevant approach was also recently introduced in [29], where the authors provide a method to measure all quadratic and quartic fermionic terms in the Jordan-Wigner encoding via quantum circuits with constant depth, assuming all-to-all qubit connectivity and noiseless multi-qubit projective measurements, with sample complexity comparable to fermionic classical shadows.
@@
In this work we consider an alternative approach for estimating non-commuting fermionic observables by implementing a simple and efficient joint measurement, in analogy to the strategy for multi-qubit Pauli observables [30]. In particular, we construct a joint measurement of a modified (noisy) collection of Majorana monomials, whose outcome statistics can be reproduced from the (parent) measurement together with an efficient classical post-processing [31, 32, 33]. Our measurement scheme for an $N$ mode fermionic system involves sampling from two distinct subsets of unitaries followed by a measurement of the occupation numbers and a simple classical post-processing. The first subset of unitaries realizes products of Majorana fermion operators. The second subset (consisting of several fermionic Gaussian unitaries) rotates disjoint blocks of the Majorana operators into balanced superpositions. For the case of quadratic and quartic monomials we show that sets of two or nine fermionic Gaussian unitaries are sufficient, respectively, to jointly measure all noisy versions of the desired observables. We also describe a tailored joint measurement scheme for estimating energies of electronic structure Hamiltonians, where four fermionic Gaussian unitaries in the second subset are sufficient. As with classical shadows, the information from a single experiment in our scheme can be used to estimate multiple observables.
@@
Our joint measurement strategy estimates expectation values of quadratic and quartic Majorana monomials to $\epsilon$ precision with $\mathcal{O}(N\log(N)/\epsilon^{2})$ and $\mathcal{O}(N^{2}\log(N)/\epsilon^{2})$ measurement rounds, respectively, and provides the same performance guarantees as fermionic (matchgate) classical shadows [1, 2]. Under the Jordan-Wigner transformation, in the setting of a rectangular lattice of qubits, the measurement circuit can be achieved with depth $\mathcal{O}(N^{1/2})$ and using $\mathcal{O}(N^{3/2})$ two-qubit gates. Fermionic and matchgate classical shadow circuits, on the other hand, require depth $\mathcal{O}(N)$ and $\mathcal{O}(N^{2})$ two-qubit gates, as well as requiring a randomization over a large collection of unitaries [1, 2]. Under standard fermion-to-qubit mappings, we estimate the expectation values of Majorana pairs and quadruples from single-qubit measurement outcomes of one and two qubits, respectively. Thus each estimate is affected only by errors on at most two qubits. As such, our strategy has the potential to be easily combined with randomized error mitigation techniques (see e.g. [34]). We benchmark our strategy with several popular molecular Hamiltonians and find that the sample complexities are comparable to other state-of-the-art schemes [24, 1, 35, 25].
@@
This work is accompanied by a complementary paper [36], which studies joint measurability of Majorana fermion operators from a more formal perspective, in particular by showing (asymptotic) optimality of our joint measurement scheme and establishing connections to graph theory, as well as topics in mathematical physics such as the SYK model.
@@
The rest of the paper is organized as follows. First, in Sec. II.1 we introduce the necessary concepts regarding fermionic systems and quantum measurements. In Sec. II.2 we present the general protocol for estimating expectation values of products of Majorana operators. In Sec. II.3 we provide the general structure of the fermionic Gaussian unitaries that guarantee a competitive scaling of the variance of the estimators for Majorana pairs and quadruples. In Sec. II.4 we specialize our scheme to quantum chemistry Hamiltonians and present its optimized implementation for a 2D qubit layout. In Sec. II.5 we present the results of numerical benchmarks of our scheme. Finally, in Sec. III we conclude the paper with a short discussion and list of open problems. In Sec. IV we summarize the methods used to obtain our results. In particular, in Sec. IV.1 we provide a proof justifying our proposed general protocol. In Sec. IV.2 we discuss the gate depth and gate count required to implement our scheme on a quantum computer under different fermion-to-qubit mappings and in different qubit layouts.
II results
II.1 Setting and notation
@@
It is convenient to introduce Majorana fermion operators $\gamma_{1},\gamma_{2},\ldots,\gamma_{2N}$, which are fermionic analogues of quadratures known from quantum optics and defined as $\gamma_{2i-1}=a_{i}+a^{\dagger}_{i}$ and $\gamma_{2i}=\mathrm{i}(a^{\dagger}_{i}-a_{i})$. Majorana operators are Hermitian and satisfy the anticommutation relations $\{{\gamma_{i},\gamma_{j}\}}=2\delta_{ij}\mathbb{I}$ In this work we are interested in estimating expectation values of (Hermitian) products of Majorana operators, with particular emphasis of monomials of degree two and four (pairs and quadruples). For even-sized subset $A\subset[2N]$ (using the convention $[k]=\{{1,\ldots,k\}}$) we define: $\gamma_{A}\coloneqq\mathrm{i}^{|A|/2}\prod_{i\in A}\gamma_{i}$, where $|A|$ denotes the size of subset $A$. We assume the convention that terms in the product are arranged in a non-decreasing manner, for example $\gamma_{1,3}=i\gamma_{1}\gamma_{3}$ and $\gamma_{1,3,6,11}=-\gamma_{1}\gamma_{3}\gamma_{6}\gamma_{11}$. For any $A,B\subset[2N]$, Majorana monomials satisfy the commutation relation $\gamma_{A}\gamma_{B}=(-1)^{|A|\cdot|B|+|A\cap B|}\gamma_{B}\gamma_{A}$, which makes it impossible to measure all observables simultaneously [37].
@@
It is often convenient to represent a fermionic system as a system of qubits via a fermion-to-qubit mapping. A transformation of this type maps the $2N$ Majorana operators to multi-qubit Pauli operators while preserving their anticommutation relations. One such mapping is the Jordan-Wigner (JW) transformation [41], in which the Majorana operators of mode $j\in[N]$ are represented as the $N$-qubit Pauli operators, $\gamma_{2j-1}=Z^{\otimes j-1}\otimes X\otimes\mathbb{I}_{2}^{\otimes N-j}$ and $\gamma_{2j}=Z^{\otimes j-1}\otimes Y\otimes\mathbb{I}_{2}^{\otimes N-j}$, where $X,Y$ and $Z$ are the qubit Pauli operators, and $\mathbb{I}_{2}$ the $2\times 2$ identity. While the Jordan-Wigner transformation can be defined irrespective of the physical arrangement of qubits, different layouts (and labellings of qubits) can have various implementation benefits (see Sec. IV.2 for further discussions and exemplary realizations of the JW transformation in 1D and 2D layouts).
@@
with $\eta_{A}\in(0,1]$ the visibility and $e_{A}\in\{\pm 1\}$ the measurement outcome (see Fig. 1 for an illustration of the scheme).
@@
of the expectation values $\operatorname{tr}(\gamma_{A}\rho)$ for an (unknown) state $\rho$. We can therefore estimate many non-commuting operators simultaneously, just as in the case of classical shadows [1, 2, 24]. The efficiency of the strategy for a set of observables $\mathcal{S}_{k}=\{\gamma_{A}:|A|=k\}$ is quantified by the sample complexity, i.e. the number of copies of the state $\rho$ needed to ensure, with probability at least $1-\delta$, that $|\operatorname{tr}(\gamma_{A}\rho)-\hat{\gamma}_{A,S}|<\epsilon$ for all $\gamma_{A}\in\mathcal{S}_{k}$, where $\hat{\gamma}_{A,S}=(1/S)\sum_{i=1}^{S}\hat{\gamma}^{(i)}_{A}$ is the empirical mean of the estimator $\hat{\gamma}_{A}$ from $S$ independent rounds. Since the estimator $\hat{\gamma}_{A}$ is binary and takes the values {±$\{\pm\eta_{A}^{-1}\}$, we can apply Hoeffding’s inequality to bound the probability for each observable [42]. Following simple reasoning (see [36]) based on the union bound (over different observables $\gamma_{A}$), we get
@@
where $|\mathcal{S}_{k}|=\binom{2n}{k}$ denotes the number of estimated observables and $\eta_{k}$ the maximal visibility for which all observables in $\mathcal{S}_{k}$ can be jointly measured.
@@
II.2 A General Joint Measurement Strategy
@@
Figure 1: Circuit implementation of the joint measurement scheme described in (i)–(iv) in Sec. II.2 for an $N$ mode fermionic state. In step (i) the state evolves under the action of a Majorana monomial $\gamma_{X}^{\dagger}$, sampled uniformly at random from the set of $2^{2N}$ Majorana monomials. In step (ii) a fermionic Gaussian unitary $U^{\dagger}$ is applied. In step (iii) the $N$ projective measurements of the disjoint Majorana pairs are performed, with outcomes $q_{1},\ldots,q_{N}$. Finally, for a given $A\subset[2N]$, one outputs $e_{A}$ according to step (iv).
@@
    Output $e_{A}=s_{AB}(-1)^{|A\cap X|}\prod_{i\in B^{\prime}}q_{i}$.
@@
For the proof see Sec. IV.1. Note that steps (i)–(iii) above are independent of the observable $\gamma_{A}$ whose noisy variant we aim to measure. In step (iv) we conduct an efficient post-processing that depends only on $A$ and $B^{\prime}$ (which can depend on $A$) and a randomly chosen subset $X$ of $[2N]$. For this reason the POVM realized in steps (i)–(iii) is a parent POVM for noisy versions of the observables $\gamma_{A}$ with visibility parameter $\eta_{A}$, for all subsets $A\subset[2N]$ of even size. The outcomes can therefore be used to estimate, simultaneously, all non-commuting observables via the estimators $\hat{\gamma}_{A}$ from Eq. (5), since $\eta_{A}^{-1}\mathbb{E}_{X,\bm{q}}[e_{A}]=\operatorname{tr}(\rho\gamma_{A})$. Next, to minimize the sample complexity we construct unitaries which aim to maximize the visibility of the measurements.
II.3 Gaussian unitaries for measurements of Majorana pairs and quadruples
Figure 2: The $202N=$ 20 Majorana operators are partitioned into 5 subsets (rows) $Y^{(r)}_{\alpha}$ (represented by light blue rounded rectangles) of cardinality 4, for each measurement setting $r=1,2$. In the first measurement setting (a), the unitary $U_{\text{sup}}$ acts via a lower flat orthogonal matrix on each row $Y_{\alpha}^{(1)}$, transforming every $\gamma_{k}$, $k\in Y_{\alpha}^{(1)}$, into a balanced superposition $\tilde{\gamma}_{k}=\sum_{l\in Y_{\alpha}^{(1)}}b^{(\alpha)}_{lk}\gamma_{l}$, with $b^{(\alpha)}_{lk}=\Omega(N^{-1/4})$. The pairings $(\pi(2j-1),\pi(2j))$ (marked by solid black lines) are constructed by permuting the standard pairings $\{(2j-1,2j)\}_{j=1}^{N}$ (represented by dark blue rounded rectangles) with a permutation $\pi$ that vertically shifts (cyclically) the second and fourth columns by 1 and 2 respectively. These pairings ensure the sample from a noisy variant of $\gamma_{ij}$ can be obtained if $i$ and $j$ are in different rows. In the second setting (b), the scheme is modified by reshuffling the operators by a permutation $\sigma^{(2)}$, that shifts columns 2,3,4 (of setting 1) vertically downwards by 1,2,3 rows, respectively. All operators in the same row of setting 1 now appear in different rows of setting 2, ensuring joint measurability of all pairs. The standard pairings represented by dark blue rectangles are consistent with the JW encoding in a 2D grid (see Fig. 4), where each such rectangle corresponds to one qubit.
To target Majorana pairs and quadruples, we propose a randomized version of the original scheme by introducing a preliminary step (o) in which every experimental run is preceded by sampling uniformly at random from a set of $K$ unitaries $U^{(r)}\in\mathrm{FLO}(N)$, $r=1,\ldots,K$. Then, we proceed with the usual steps (i)–(iii), as described in Sec. II.2, with $U^{(r)}$ implemented in (iii). Here we sketch the construction of the $K$ unitaries that ensures that such a modification of the joint measurement from Prop. 1 gives visibilities $\eta_{2}=\Theta(1/\sqrt{N})$ and $\eta_{4}=\Theta(1/N)$, where $\eta_{k}$ denotes the maximum visibility for which all $k$–th degree Majorana monomials can be jointly measured by the parent. The details of the construction together with an extension to arbitrary constant-sized Majorana monomials are presented in [36].